OpenStudy (anonymous):

1.7t + 8 -1.62 = 0.4t - 0.32t +8

OpenStudy (anonymous):

1.7t + 8 - 1.62=0.4t - 0.32t + 8 1.7t - 0.4t + 0.32t = 8 -8 + 1.62 1.62t = 1.62 t=1

OpenStudy (anonymous):

thanks